Please help me to solve
Use the elimination method to solve the system.
2x+2y=-14
2x-3y=-4

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Multiply the 1st eqn by 3 and the 2nd by 2, then add them.

Answer by ewatrrr(24785)

 
Hi,
Use the elimination method to solve the system.
2x+2y=-14
2x-3y=-4
2x+2y=-14
-2x+3y=+4 | Multiplying 2nd EQ thru by -1 to eliminate the x variable
5y = -10
y = -2 and x = -5 | 2x = -4 + 3*(-2), x = -10/2
CHECKING our Answer***
2*-5 + 2*-2 = -14